# Windowed MAPF with Completeness Guarantees

### v1

@@ C.1 Subtleties
Technically, given a set of possible heuristic penalties, we could try to maximize the combination of these penalties such that chosen penalties do not have overlapping agents. This ends up being a maximum weighted disjoint set-cover problem where each heuristic penalties is a set and the set’s weight is the penalty. We note that solving just disjoint set-cover is NP-Complete (Cardei and Du 2005), so solving this efficiently is hard. We thus instead do a greedy approximation and choose to apply the heuristic conflict with the highest penalty.
@@
Thus, every time we compute heuristic penalties and conflicts, we recompute it from scratch while maintaining consistency with prior vertex constraints. This allows us to initially pick $(R_{1},R_{2})$ and then afterwards “reassign” $(R_{1},R_{2},R_{3})$ later on.
Tie-Breaking
Given the 1-step plans, it is likely that there are many SS-CBS solutions with an equally good cost. We found it helpful to tiebreak by introducing agent “priorities”. The intuition behind this is that given an symmetric situation of two agents in a hallway (where all adjacent configurations have the same summed $h$ value), we prefer one agent to “push” the other agent away rather than oscillate in the middle.
Concretely, if two CT nodes have the same f-value, h-value, g-value, and conflict, we tie-break nodes by lexicographically comparing the agent’s h-values. This means that given equally good options, we would prefer a solution that reduces the first lexographically sorted agent’s heuristic more than other options. The agent’s lexicographical ordering can be thought of as it’s tiebreaking priority (i.e. the agent that comes first is the highest priority). We found assigning random “priorities” that get updated as in PIBT to noticeably help performance compared to not having it (see ablations for more details).
C.2 SS-CBS Small Congested Maps
Table A1 shows the success rate and average runtime in milliseconds for three extremely tough MAPF maps with number of agents in parenthesis. We run 20 seeds with a timeout of 1 minute, with the number $N$ in parenthesis denoting using the first $N$ agents in a scenario. The time is over successful instances in milliseconds. We ran windowed ECBS with window sizes $1,2,4,8,16$ but these baselines all failed (as shown in Table 1). Thus we compare it to EECBS with a full planning horizon and suboptimality $w=1,1.5$. We highlight that we use EECBS (Explicit Estimation high-level search) with all CBS improvements including bypass, symmetry reasoning, prioritized conflicts, and full horizon planning. SS-CBS uses no CBS improvements with the basic focal search with $w_{so}=1$ and plans a single step.
We see how SS-CBS is able to have a consistently higher success rate and lower runtime compared to EECBS. SS-CBS’s performance is particularly impressive in Loopchain where EECBS is completely unable to solve this scenario. Additionally, on solved instances, SS-CBS has runtime faster by an order of magnitude compared to EECBS.
SS-CBS EECBS $w_{so}=1$ EECBS $w_{so}=1.5$
Instance Success Time Success Time Success Time
Tunnel (3) 1 85 0.9 43423 0.8 3768
Tunnel (4) 1 2713 0 - 0 -
Loopchain (6) 1 15393 0 - 0 -
Loopchain (7) 0.95 18268 0 - 0 -
Connector (5) 1 15 0 - 0.95 146
Connector (6) 1 43 0 - 0.9 1116
Table A1: SS-CBS with one-step planning can solve problems that even EECBS with full horizon planning and all optimizations cannot. Time in milliseconds, timeout of 1 minute. This highlights SS-CBS’s and the broader WinC-MAPF framework’s ability to effectively use grouped heuristic penalties to resolve congestion.
@@ C.3 SS-CBS Ablation Study
Figure A1 shows an ablation study of SS-CBS by varying the high-level suboptimality and tiebreaking mechanism. We test suboptimalities of $w_{so}=1,1.05,1.1,1.5$ as well as explore two tie-breaking variants along with the none tie-breaking base version. The none tie-breaking version will arbitrarily pick any node that shares the same f, h, g, and conflict value. We note that the tie-breaking with optimal SS-CBS (e.g. $w_{so}=1$) is theoretically complete under our proof. Using $w_{so}>1$ is not theoretically proven or disproven to be complete and is left for future work to study its theoretical properties.
@@
In Figure A1, colors denote different high-level suboptimalities while the linestyle denotes different tie-breaks. The “-n” corresponds to no tie-breaking scheme, “-d” to an initial Distance based priority, and “-r” to an initial random based priority. The first row shows how increasing the suboptimality can improve the scalability of SS-CBS. However interestingly we see that except for random-32-32-20, $w_{so}=1.05,1.1,1.5$ have near identical performance. The first row also highlights how including the tie-breaking mechanisms has a small but noticeable improvement over no tie-breaks. A close inspection shows that the random initialization tie-breaking helps more than the pibt initialization tie-break.
The second row shows the number of iterations of execution require to find the goal (equivalent to makespan). We see two possible patterns given a problem instance where multiple suboptimalities find a solution. In random-32-32-20 or warehouse, we generally see that increasing the suboptimality increases the number of iterations required compared to lower suboptimalities. Upon visual inspection, this seemed to occur when there were instances of deadlock that needed to be resolved via HPs. Higher suboptimalities meant that instead of trying to navigate outside of encountered HPs, SS-CBS several times chose to pick the same location and incur the higher HP cost. However in the less congested maps ht_chantry and den520d, we do not see as strong a pattern. Thus higher suboptimality seems to incur this drawback mainly in congested scenarios. We see again that tie-breaking helps reduce the number of iterations required to solve problems.
The last two row shows how many HPs were created and how many HP conflicts were found in CTs across all search iterations. We see that the number of HPs increases roughly linearly as the number of agents increase. However, the number of HPs found increases exponentially as the number of agents increases. This highlights how SS-CBS spends more effort on resolving congestion via HPs as the number of agents (and therefore congestion) increases.
@@
We see that on 3 of the 4 maps, increasing the suboptimality from 1 to 1.05 leads to an improvement but that higher suboptimalities does not for SS-CBS. We visually observed that SS-CBS would would be fine waiting at a location with HP when the suboptimality increases (as the suboptimality factor allowed that solution). This shows how more clever ways of incorporating suboptimal search and real-time searches must be researched, similar to Weighted-LRTA* (Rivera, Baier, and Hernandez 2013) which showed a non-trivial adjustment to LRTA* to work with sub-optimal searches.
@@
Figure A1: We plot statistics of SS-CBS with different high-level suboptimality (colored) and CT tie-breaking (line style).

### v2

@@ C.1 Determining Which HPs to Apply
Given a set of possible HPs, we could try to maximize the combination of these penalties such that chosen penalties do not have overlapping agents. This ends up being a maximum weighted disjoint set-cover problem where each heuristic penalties is a set and the set’s weight is the penalty. We note that solving just disjoint set-cover is NP-Complete (Cardei and Du 2005), so solving this efficiently is hard. We thus instead do a greedy approximation and choose to apply the heuristic conflict with the highest penalty.
@@
Thus, every time we compute heuristic penalties and conflicts, we recompute it from scratch while not violating constraints. This allows us to initially pick $(R_{1},R_{2})$ and then afterwards “reassign” $(R_{1},R_{2},R_{3})$ later on.
C.2 Tie-Breaking
Given the 1-step plans, it is likely that there are many SS-CBS solutions with an equally good cost. We found it helpful to tiebreak by introducing agent “priorities”. The intuition is that given a symmetric situation of two agents in a hallway (where all adjacent configurations have the same summed $h$ value), we prefer one agent to “push” the other agent away rather than oscillate in the middle.
Concretely, if two CT nodes have the same f-value, h-value, g-value (cost), and number of conflicts, we tie-break nodes by lexicographically comparing the agent’s h-values. This means that given equally good options, we would prefer a solution that reduces the first lexicographically sorted agent’s heuristic more than other options. The agent’s lexicographical ordering can be thought of as its tiebreaking priority (i.e., the agent that comes first is the highest priority). We additionally found assigning random “priorities” which are updated as in PIBT helped performance.
@@ C.3 SS-CBS Ablation Study
A powerful technique for speeding up CBS is Enhanced CBS (Barer et al. 2014) which replaces the low-level and high-level optimal searches with bounded-suboptimal focal searches. Thus, an obvious idea to potentially improve SS-CBS is to introduce these focal searches. As stated in Section 6, our proof of completeness only applies to optimal AGs. For experimental curiosity, we tested out SS-CBS with a suboptimal high-level search (although we do not prove completeness for this suboptimal AG).
Figure A1 shows an ablation study of SS-CBS by varying the high-level suboptimality and tiebreaking mechanism. We test suboptimalities of $w_{so}=1,1.05,1.1,1.5$ as well as explore two tie-breaking variants along with the none tie-breaking base version. The none tie-breaking version will arbitrarily pick any node that shares the same f, h, g, and number of conflicts. Using $w_{so}>1$ is not theoretically proven or disproven to be complete and is left for future work to study its theoretical properties.
@@
In Figure A1, colors denote different high-level suboptimalities while the linestyle denotes different tie-breaks. The “-n” corresponds to no tie-breaking scheme, “-d” to an initial Distance based priority, and “-r” to an initial random based priority. The first row shows how increasing the suboptimality can improve the scalability of SS-CBS. The first row also highlights how including the tie-breaking mechanisms has a small but noticeable improvement over no tie-breaks. A close inspection shows that the random initialization tie-breaking helps more than the PIBT initialization tie-break.
The second row shows the number of iterations of execution required to reach the goal (equivalent to makespan). We see two possible patterns given a problem instance where multiple suboptimalities find a solution. In random-32-32-20 or warehouse, we generally see that increasing the suboptimality increases the number of iterations required compared to lower suboptimalities. Upon visual inspection, this seemed to occur when there were instances of deadlock that needed to be resolved via HPs. Higher suboptimalities meant that instead of trying to navigate outside of encountered HPs, SS-CBS several times chose to pick the same location and incur the higher HP cost. However, in the less congested maps ht_chantry and den520d, we do not see as strong a pattern. Thus higher suboptimality seems to incur this drawback mainly in congested scenarios. We see again that tie-breaking helps reduce the number of iterations required to solve problems.
The last two rows show how many HPs were created and how many HP conflicts were found in CTs across all search iterations. We see the number of HPs found in search increases exponentially as the number of agents increases, although the total number of HPs created does not increase as much. This highlights how SS-CBS spends more effort on resolving congestion via HPs as the number of agents (and therefore congestion) increases.
@@
We see that on 3 of the 4 maps, increasing the suboptimality from 1 to 1.05 leads to an improvement but that higher suboptimalities does not for SS-CBS. We visually observed that SS-CBS would be fine waiting at a location with HP when the suboptimality increases (as the suboptimality factor allowed that solution). This shows how more clever ways of incorporating suboptimal search and real-time searches must be researched, similar to Weighted-LRTA* (Rivera, Baier, and Hernandez 2013) which showed a non-trivial adjustment to LRTA* to work with sub-optimal searches.
@@
Figure A1: We plot statistics of SS-CBS with different high-level suboptimality $w_{so}$ (colored) and CT tie-breaking (line style). Note that completeness is only proven for an optimal SS-CBS (i.e., $w_{so}=1$ and not $w_{so}>1$).